OpenStudy (anonymous):

In the graph of a power function y=kx^p, what form does the graph take if p is even?

OpenStudy (anonymous):

I'm sorry? What comment? Also I thought parabolas where just when the exponent was 2

OpenStudy (loser66):

for even exponents, all are parabolas

OpenStudy (e.mccormick):

Some examples https://www.desmos.com/calculator/yyun6m8vhc

OpenStudy (anonymous):

This is exactly what I was looking for, thank you e.mc

OpenStudy (e.mccormick):

It is basically what L66 said, A bunch of U shapes, parabolas, etc. They might be wider or narrower, even upside down, but all have two tails going in the same direction to \(\pm \infty\). As opposed to all odd powers, which have one going to \(+\infty\) and one going to \(-\infty\).

OpenStudy (anonymous):

Yeah I was just confused because I saw no choice like that, I had forgotten that the exponent could also be odd haha